One kg of a steam with quality of 20 percent is heated at a constant pressure of 200kPa until the temperature reaches 400°C. Calculate the work done by the system.
1
Expert's answer
2019-06-11T10:01:52-0400
"W=mP(v_2-v_1)"

From steam tables:

"v_1=0.8857\frac{m^3}{kg}"

"v_2=1.5493\frac{m^3}{kg}"

"W=(1)(200000)(1.5493-0.8857)=132700\ J=132.7\ kJ"